(x^2-2x)(x+3)=neg.2x(x+1)
Answer by stanbon(75887)

(x^2-2x)(x+3)=neg.2x(x+1)
------
x^3 + 3x^2 -2x^2 -6x = -2x^2 -2x
----
x^3 + 3x^2 -4x = 0
x(x^2 + 3x - 4) = 0
x(x+4)(x-1) = 0
x = 0 or x = -4 or x = 1
============================
Cheers,
Stan H.